OpenStudy (anonymous):

A marketing company found that 4% of people who were going to buy a new television wanted the largest screen possible. If 1600 people were going to buy a television next year, how many would be expected to want to buy the largest screen possible?

OpenStudy (anonymous):

6400 or 400?

OpenStudy (anonymous):

\[\frac{ 4 }{ 100 }=0.04*1600=answer\]

OpenStudy (anonymous):

._. ok thanks

OpenStudy (anonymous):

No Problem, hope that helps :)

OpenStudy (anonymous):

so 64?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thanks